OpenStudy (anonymous):

Super hard math Problem! A wagon train that is one mile long advances one mile at a constant rate. During the same time period, the wagon master rides his horse at a constant rate from the front of the wagon train to the rear, and then back to the front. How far did that wagon master ride?

OpenStudy (anonymous):

let train reached a fraction of it's journey, t, when horse arrived at train's rear (covering 1-t of the train's length). so when train go on with the rest of it's journey, 1-t, the horse had to cover 2-t in the same time frame. letting c be train's speed and v for horse's speed, v / c = (1 - t) / t = (2 - t) / (1 - t) (1 - t)^2 = t(2 - t) t^2 - 2t + 1 = 2t - t^2 2t^2 - 4t + 1 = 0 2(t - 1)^2 = 1 t = 1 +- √0.5 since 1 - t > 0, t = 1 - √0.5 v / c = (2 - t) / (1 - t) = (1 + √0.5) / √0.5 = √2 + 1 distance travelled by horse (and the rider, wagon master) = (√2 + 1) * 1 = 2.41421 miles

OpenStudy (anonymous):

OMGGGG, you saved my life! Thank you so much for your help!!!